# Deep Learning and Machine Learning - Python Data Structures and Mathematics Fundamental: From Theory to Practice

### v1

@@ 25.2.2 Simpson’s Rule
Simpson’s rule provides a more accurate approximation of integrals by using quadratic polynomials to approximate the function within each subinterval[215, 29]. The formula for Simpson’s rule is:
[…]
@@ 25.2.3 Gaussian Quadrature
Gaussian quadrature is a powerful technique for numerical integration that provides exact results for polynomials of degree $12n-$ 1 or less, where $n$ is the number of sample points[96]. It selects both the sample points and weights optimally to achieve high accuracy.
@@
In this example, we used Gaussian quadrature to approximate the integral of $\sin(x)$ over $[0,\pi]$. Gaussian quadrature is particularly useful for high-precision integration.
@@ 26.1 Direct Methods
Direct methods aim to solve a system of linear equations in a finite number of steps, usually through matrix factorizations[70, 238, 120, 269, 107]. These methods are precise but can be computationally expensive for large matrices. Common direct methods include Gaussian elimination, LU decomposition, and Cholesky decomposition.
@@ 26.1.1 Gaussian Elimination
Gaussian elimination is a method for solving linear systems by converting the system’s matrix into an upper triangular form[159, 260, 220, 107, 269]. Once the matrix is in this form, the solution can be obtained through back-substitution.
@@ 26.1.2 LU Decomposition
LU decomposition is a method that factors a matrix $A$ into the product of two matrices: a lower triangular matrix $L$ and an upper triangular matrix $U$[288, 77, 107, 259]. This is useful for solving linear systems because once $A$ is decomposed, solving the system becomes a matter of solving two triangular systems.
@@ 26.1.3 Cholesky Decomposition
Cholesky decomposition is a specialized version of LU decomposition for symmetric, positive-definite matrices[53, 215, 269, 106]. It decomposes a matrix $A$ into the product of a lower triangular matrix $L$ and its transpose:
@@ 26.2.1 Jacobi Method
The Jacobi method is an iterative algorithm for solving linear systems. It updates each variable in the system independently of the others using the previous iteration’s values[8, 238, 297, 275, 21]. The system $A\mathbf{x}=\mathbf{b}$ is written as:
@@ 26.2.2 Gauss-Seidel Method
The Gauss-Seidel method improves on the Jacobi method by using updated values as soon as they are available in the iteration[238, 71, 107, 269, 37]. This makes the Gauss-Seidel method faster than the Jacobi method for many problems.
@@ 26.2.3 Conjugate Gradient Method
The Conjugate Gradient method is an efficient iterative algorithm for solving large, sparse systems of linear equations, especially when the matrix is symmetric and positive-definite[119, 85, 200, 107]. The method seeks to minimize a quadratic form iteratively.
@@ 27.1 Matrix Factorization
Matrix factorization is a fundamental tool in numerical linear algebra[151, 103]. It refers to the process of decomposing a matrix into a product of matrices with certain properties. The most common types of matrix factorization used in machine learning and deep learning include Singular Value Decomposition (SVD) [79, 107, 105]and QR Decomposition[125, 122, 110].
@@ 27.1.1 Singular Value Decomposition (SVD)
[…]
@@ 27.2 Eigenvalues and Eigenvectors
Eigenvalues and eigenvectors are fundamental concepts in linear algebra with wide-ranging applications in machine learning, physics, and data analysis[265, 24]. For a given square matrix $A$, an eigenvector $\mathbf{v}$ and its corresponding eigenvalue $\lambda$ satisfy the equation:
@@ 27.3 Principal Component Analysis (PCA)
Principal Component Analysis (PCA) is a statistical technique used for dimensionality reduction[213, 124, 91, 137, 116]. It is based on finding the directions (principal components) in which the data varies the most. PCA transforms the original high-dimensional data into a lower-dimensional space, while preserving as much variability as possible.
@@ 28.1 Introduction to Fourier Transform
The Fourier Transform decomposes a function into its constituent frequencies[88, 26, 253]. It transforms a signal from the time domain, where the signal is expressed as a function of time, to the frequency domain, where the signal is expressed in terms of its frequency components.
@@ 28.1.1 Mathematical Definition of the Fourier Transform
The continuous Fourier Transform (FT)[253, 27] of a function $f(t)$ is defined as:
[…]
@@
The inverse Fourier Transform allows us to reconstruct the original function from its frequency components[286, 73]:
[…]
@@ 28.2 Discrete Fourier Transform (DFT)
The Fourier Transform for continuous signals assumes that the signal is sampled at infinite points, but in practice, we deal with discrete data, such as digital signals. The Discrete Fourier Transform (DFT) is used to analyze finite, discrete sequences[59, 224, 205].
@@ 28.2.1 Mathematical Definition of the DFT
  • $N$ is the total number of points.
@@ 28.3 Fast Fourier Transform (FFT)
The Fast Fourier Transform (FFT) is an efficient algorithm for computing the DFT, reducing the time complexity from $O(N^{2})$ to $O(N\log N)$. FFT is one of the most important algorithms in numerical computing because it allows the analysis of large datasets quickly[59, 38].
@@ 28.4.3 Convolution Theorem in Deep Learning
In deep learning, convolutions are a fundamental operation in convolutional neural networks (CNNs). The Convolution Theorem states that convolution in the time (or spatial) domain is equivalent to multiplication in the frequency domain[209, 205, 26, 236]. This can significantly speed up the convolution operation, especially for large images or 3D data.
@@
In this chapter, we will introduce nonlinear systems, explain widely used methods such as Newton’s method [221] and Broyden’s method[31, 40] for solving nonlinear systems[207], and explore their applications in optimization tasks in neural networks.
@@ 29.2 Newton’s Method for Nonlinear Systems
Newton’s method is one of the most popular iterative methods for solving systems of nonlinear equations. It extends the basic idea of Newton’s method for scalar functions to systems of equations. The method relies on the Jacobian matrix, which contains the partial derivatives of each equation with respect to each variable. The Jacobian matrix is used to iteratively improve an initial guess until the solution converges[162, 250].
@@
The Jacobian matrix $J_{F}$ for a system of equations is defined as[94, 12]:
@@ 29.3 Broyden’s Method
Broyden’s method is a quasi-Newton method for solving systems of nonlinear equations[172, 309, 201]. While Newton’s method requires the computation of the Jacobian matrix at each iteration, Broyden’s method updates an approximation to the Jacobian, reducing computational cost. This makes Broyden’s method useful when the Jacobian is expensive to compute or when it is not readily available[289, 200, 107, 262].
@@ 29.3.1 Broyden’s Method Algorithm
Broyden’s method starts with an initial guess for the solution $x_{0}$ and an initial approximation to the Jacobian $B_{0}$. The algorithm proceeds iteratively[212]:
@@ 29.4 Applications in Optimization for Neural Networks
The optimization process involves finding the parameter vector $\theta$ that minimizes the loss function $L(\theta)$[246]. Gradient-based methods such as gradient descent or more advanced techniques like quasi-Newton methods (e.g., Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm[32]) are commonly used to solve this optimization problem.
@@ 20 H = hessian(theta)
Numerical optimization refers to the process of finding the minimum or maximum of a function when an analytical solution is difficult or impossible to obtain. Optimization is fundamental in many fields such as machine learning, physics, economics, and engineering[100, 86, 24]. In this chapter, we will explore various numerical optimization techniques, starting from gradient-based methods and advancing to more sophisticated approaches like quasi-Newton methods and gradient-free methods. We will also explore how these methods are applied in training deep neural networks.
@@ 30.1.1 Types of Optimization Problems
  • Unconstrained Optimization[39, 291, 58, 242, 184, 118]: In this case, there are no restrictions on the values that $x$ can take. We aim to find the global or local minimum of the objective function.
  • Constrained Optimization[142, 218, 249, 163, 111, 183, 214]: Here, the variable $x$ is subject to certain constraints, such as $g(x)\leq 0$ or $h(x)=0$. The optimization needs to account for these constraints.
@@ 30.2 Gradient-Based Methods
Gradient-based methods use the gradient (first derivative) of the objective function to guide the optimization process[83]. These methods are particularly efficient for smooth, differentiable functions[153]. The key idea is that the gradient indicates the direction of the steepest ascent or descent of the function.
@@ 30.2.2 Conjugate Gradient Method
The conjugate gradient method is a more efficient alternative to gradient descent, especially for large-scale optimization problems where the objective function is quadratic or approximately quadratic[300, 283]. It minimizes the function along conjugate directions rather than along the gradient direction alone[119].
@@ 30.3 Quasi-Newton Methods
Quasi-Newton methods are second-order optimization techniques that approximate the Hessian matrix (the matrix of second derivatives) to improve convergence[171, 146, 281, 263]. These methods are computationally more efficient than full Newton’s method, which requires the computation of the full Hessian matrix[304].
@@ 30.3.2 L-BFGS Algorithm
The Limited-memory BFGS (L-BFGS) algorithm is a memory-efficient version of BFGS[170, 251, 95]. Instead of storing the full inverse Hessian matrix, it stores only a few vectors to approximate it. This makes L-BFGS suitable for large-scale problems with many variables[164, 307].
@@ 30.4 Gradient-Free Optimization
Gradient-free optimization methods are useful when the objective function is not differentiable, noisy, or when computing the gradient is computationally expensive[197, 51, 305]. These methods do not rely on gradient information and instead explore the search space based on the function values[138].
@@ 30.4.1 Nelder-Mead Method
The Nelder-Mead method is a popular gradient-free optimization algorithm. It uses a simplex of $n+1$ points in $n$-dimensional space and iteratively adjusts the simplex to converge to the minimum of the function[161]. The algorithm involves operations like reflection, expansion, contraction, and shrinking to adjust the simplex[230].
@@ 30.5 Applications in Training Deep Neural Networks
SGD is a variant of gradient descent that updates the parameters using a small subset (mini-batch) of the training data at each step[232, 23]. This makes it more efficient for large datasets. The update rule for SGD is:
[…]
@@
6class SimpleNet(nn.Module):
@@
Ordinary Differential Equations (ODEs) are equations that describe the relationship between a function and its derivatives[6, 30]. They play a crucial role in many fields of science and engineering, including neural networks and deep learning, where they are used to model dynamic systems[48, 235]. In this chapter, we will cover the basic concepts of ODEs, methods for solving them, and their applications in modeling neural dynamics.
@@ 31.2 Euler’s Method
Euler’s method is the simplest numerical technique for solving ODEs[147]. It approximates the solution of an ODE by taking small steps along the curve of the solution. Given the ODE:
@@
  • We define the function $f(t,y)=-2y$ and solve it using Euler’s method.
@@ 31.3 Runge-Kutta Methods
Euler’s method, while simple, is not very accurate for small step sizes. Runge-Kutta methods are more advanced numerical techniques that provide better accuracy for solving ODEs[237, 75]. The most commonly used method is the fourth-order Runge-Kutta method (RK4), which provides a much better approximation of the solution by using a weighted average of several slopes[154, 182, 135].
@@ 31.4 Stiff ODEs
Stiff ODEs are a special class of differential equations where certain numerical methods (such as Euler’s method) can become unstable unless very small step sizes are used[97, 113, 114]. Stiffness typically arises in systems where there are processes that occur on vastly different time scales[82, 293, 46, 186, 268].

### v2

@@ 25.2.2 Simpson’s Rule
Simpson’s rule provides a more accurate approximation of integrals by using quadratic polynomials to approximate the function within each subinterval[phillips1962new, brown1967introduction]. The formula for Simpson’s rule is:
$\int_{a}^{b}f(x)\,dx\approx\frac{h}{3}\left[f(a)+4\sum_{i=1,3,5,\dots}^{n-1}f(x_{i})+2\sum_{i=2,4,6,\dots}^{n-2}f(x_{i})+f(b)\right]$
@@ 25.2.3 Gaussian Quadrature
Gaussian quadrature is a powerful technique for numerical integration that provides exact results for polynomials of degree 2n12n-1 or less, where $n$ is the number of sample points[gauss1814legendre]. It selects both the sample points and weights optimally to achieve high accuracy.
@@
In this example, we used Gaussian quadrature to approximate the integral of $\sin(x)$ over $0,\pi]$. Gaussian quadrature is particularly useful for high-precision integration.
@@ 26.1 Direct Methods
Direct methods aim to solve a system of linear equations in a finite number of steps, usually through matrix factorizations[demmel1989stability, saad2003iterative, higham1990stable, trefethen1997numerical, golub2013matrix]. These methods are precise but can be computationally expensive for large matrices. Common direct methods include Gaussian elimination, LU decomposition, and Cholesky decomposition.
@@ 26.1.1 Gaussian Elimination
Gaussian elimination is a method for solving linear systems by converting the system’s matrix into an upper triangular form[lay2015linear, strang2016introduction, pratt1977stable, golub2013matrix, trefethen1997numerical]. Once the matrix is in this form, the solution can be obtained through back-substitution.
@@ 26.1.2 LU Decomposition
LU decomposition is a method that factors a matrix $A$ into the product of two matrices: a lower triangular matrix $L$ and an upper triangular matrix UU[wilkinson1965rounding, duff1983ma48, golub2013matrix, strang1988linear]. This is useful for solving linear systems because once $A$ is decomposed, solving the system becomes a matter of solving two triangular systems.
@@ 26.1.3 Cholesky Decomposition
Cholesky decomposition is a specialized version of LU decomposition for symmetric, positive-definite matrices[chol:cholesky1907, phillips1962new, trefethen1997numerical, golub2012matrix]. It decomposes a matrix $A$ into the product of a lower triangular matrix $L$ and its transpose:
@@ 26.2.1 Jacobi Method
The Jacobi method is an iterative algorithm for solving linear systems. It updates each variable in the system independently of the others using the previous iteration’s values[axelsson1996iterative, saad2003iterative, young1971iterative, varga2000matrix, bjorck1996numerical]. The system $A\mathbf{x}=\mathbf{b}$ is written as:
@@ 26.2.2 Gauss-Seidel Method
The Gauss-Seidel method improves on the Jacobi method by using updated values as soon as they are available in the iteration[saad2003iterative, demmel1989qr, golub2013matrix, trefethen1997numerical, burden2016numerical]. This makes the Gauss-Seidel method faster than the Jacobi method for many problems.
@@ 26.2.3 Conjugate Gradient Method
The Conjugate Gradient method is an efficient iterative algorithm for solving large, sparse systems of linear equations, especially when the matrix is symmetric and positive-definite[hestenes1952methods, fletcher1970new, nocedal2006numerical, golub2013matrix]. The method seeks to minimize a quadratic form iteratively.
@@ 27.1 Matrix Factorization
Matrix factorization is a fundamental tool in numerical linear algebra[koren2009matrix, golub1965computing]. It refers to the process of decomposing a matrix into a product of matrices with certain properties. The most common types of matrix factorization used in machine learning and deep learning include Singular Value Decomposition (SVD) [eckart1936approximation, golub2013matrix, golub1970singular]and QR Decomposition[householder1964methods, horn2012matrix, gu1994qr].
@@ 27.1.1 Singular Value Decomposition (SVD)
10print("Matrix $\n", U$)
11print("Singular values ($\n", S)$
12print("Matrix $\n", VT$)
@@ 27.2 Eigenvalues and Eigenvectors
Eigenvalues and eigenvectors are fundamental concepts in linear algebra with wide-ranging applications in machine learning, physics, and data analysis[tao2012topics, boyd2004convex]. For a given square matrix $A$, an eigenvector $\mathbf{v}$ and its corresponding eigenvalue $\lambda$ satisfy the equation:
@@ 27.3 Principal Component Analysis (PCA)
Principal Component Analysis (PCA) is a statistical technique used for dimensionality reduction[pearson1901liii, hotelling1933analysis, friedman1974projection, jolliffe2002principal, hastie2015statistical]. It is based on finding the directions (principal components) in which the data varies the most. PCA transforms the original high-dimensional data into a lower-dimensional space, while preserving as much variability as possible.
@@ 28.1 Introduction to Fourier Transform
The Fourier Transform decomposes a function into its constituent frequencies[fourier1822analytical, bracewell1986fourier, stein2003fourier]. It transforms a signal from the time domain, where the signal is expressed as a function of time, to the frequency domain, where the signal is expressed in terms of its frequency components.
@@ 28.1.1 Mathematical Definition of the Fourier Transform
The continuous Fourier Transform (FT)[stein2003fourier, bracewell2000fourier] of a function $f(t)$ is defined as:
$F(\omega)=\int_{-\infty}^{\infty}f(t)e^{-i\omega t}\,dt$
@@
The inverse Fourier Transform allows us to reconstruct the original function from its frequency components[watson1995treatise, dirichlet1829convergence]:
$f(t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}F(\omega)e^{i\omega t}\,d\omega$
@@ 28.2 Discrete Fourier Transform (DFT)
The Fourier Transform for continuous signals assumes that the signal is sampled at infinite points, but in practice, we deal with discrete data, such as digital signals. The Discrete Fourier Transform (DFT) is used to analyze finite, discrete sequences[cooley1965algorithm, rabiner1975theory, oppenheim1975digital].
@@ 28.2.1 Mathematical Definition of the DFT
  • NN is the total number of points.
@@ 28.3 Fast Fourier Transform (FFT)
The Fast Fourier Transform (FFT) is an efficient algorithm for computing the DFT, reducing the time complexity from $O(N^{2})$ to $O(N\log N)$. FFT is one of the most important algorithms in numerical computing because it allows the analysis of large datasets quickly[cooley1965algorithm, burrus1985fft].
@@ 28.4.3 Convolution Theorem in Deep Learning
In deep learning, convolutions are a fundamental operation in convolutional neural networks (CNNs). The Convolution Theorem states that convolution in the time (or spatial) domain is equivalent to multiplication in the frequency domain[papoulis1962fourier, oppenheim1975digital, bracewell1986fourier, rudin1962fourier]. This can significantly speed up the convolution operation, especially for large images or 3D data.
@@
In this chapter, we will introduce nonlinear systems, explain widely used methods such as Newton’s method [press1992numerical] and Broyden’s method[broyden1965class, cao2023solving] for solving nonlinear systems[ortega2000iterative], and explore their applications in optimization tasks in neural networks.
@@ 29.2 Newton’s Method for Nonlinear Systems
Newton’s method is one of the most popular iterative methods for solving systems of nonlinear equations. It extends the basic idea of Newton’s method for scalar functions to systems of equations. The method relies on the Jacobian matrix, which contains the partial derivatives of each equation with respect to each variable. The Jacobian matrix is used to iteratively improve an initial guess until the solution converges[lee2023jacobian, smith2023jacobian].
@@
The Jacobian matrix $J_{F}$ for a system of equations is defined as[garcia2023jacobian, baker2024jacobian]:
@@ 29.3 Broyden’s Method
Broyden’s method is a quasi-Newton method for solving systems of nonlinear equations[liu2023comprehensive, zhou2024quasi, nocedal2023recent]. While Newton’s method requires the computation of the Jacobian matrix at each iteration, Broyden’s method updates an approximation to the Jacobian, reducing computational cost. This makes Broyden’s method useful when the Jacobian is expensive to compute or when it is not readily available[wong2023jacobian, nocedal2006numerical, golub2013matrix, strogatz2014nonlinear].
@@ 29.3.1 Broyden’s Method Algorithm
Broyden’s method starts with an initial guess for the solution $x_{0}$ and an initial approximation to the Jacobian $B_{0}$. The algorithm proceeds iteratively[patel2024trends]:
@@ 29.4 Applications in Optimization for Neural Networks
The optimization process involves finding the parameter vector $\theta$ that minimizes the loss function $L(\theta)$[shanno1970conditioning]. Gradient-based methods such as gradient descent or more advanced techniques like quasi-Newton methods (e.g., Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm[broyden1970convergence]) are commonly used to solve this optimization problem.
@@ 20 H = hessian(theta)
Numerical optimization refers to the process of finding the minimum or maximum of a function when an analytical solution is difficult or impossible to obtain. Optimization is fundamental in many fields such as machine learning, physics, economics, and engineering[gill1981practical, fletcher2013practical, boyd2004convex]. In this chapter, we will explore various numerical optimization techniques, starting from gradient-based methods and advancing to more sophisticated approaches like quasi-Newton methods and gradient-free methods. We will also explore how these methods are applied in training deep neural networks.
@@ 30.1.1 Types of Optimization Problems
  • Unconstrained Optimization[byrd2023recent, wright2024efficient, conn2023advances, schmidt2024machine, martinez2023trust, hazra2024review]: In this case, there are no restrictions on the values that $x$ can take. We aim to find the global or local minimum of the objective function.
  • Constrained Optimization[karush1939minima, powell1978algorithms, smith2023recent, lee2023ai, gupta2023constrained, martinez2024new, perez2023constrained]: Here, the variable $x$ is subject to certain constraints, such as $g(x)\leq 0$ or $h(x)=0$. The optimization needs to account for these constraints.
@@ 30.2 Gradient-Based Methods
Gradient-based methods use the gradient (first derivative) of the objective function to guide the optimization process[fiacco1968nonlinear]. These methods are particularly efficient for smooth, differentiable functions[kuhn1951nonlinear]. The key idea is that the gradient indicates the direction of the steepest ascent or descent of the function.
@@ 30.2.2 Conjugate Gradient Method
The conjugate gradient method is a more efficient alternative to gradient descent, especially for large-scale optimization problems where the objective function is quadratic or approximately quadratic[zhang2024engineering, wang2023hybrid]. It minimizes the function along conjugate directions rather than along the gradient direction alone[hestenes1952methods].
@@ 30.3 Quasi-Newton Methods
Quasi-Newton methods are second-order optimization techniques that approximate the Hessian matrix (the matrix of second derivatives) to improve convergence[liu2023stability, kim2024high, wang2023hessian, sun2023hessian]. These methods are computationally more efficient than full Newton’s method, which requires the computation of the full Hessian matrix[zhang2023large].
@@ 30.3.2 L-BFGS Algorithm
The Limited-memory BFGS (L-BFGS) algorithm is a memory-efficient version of BFGS[liu1989limited, smith2023advances, garcia2024efficient]. Instead of storing the full inverse Hessian matrix, it stores only a few vectors to approximate it. This makes L-BFGS suitable for large-scale problems with many variables[lee2023survey, zhang2023modified].
@@ 30.4 Gradient-Free Optimization
Gradient-free optimization methods are useful when the objective function is not differentiable, noisy, or when computing the gradient is computationally expensive[nelder1965simplex, chen2024gradientfree, zhang2023evolutionary]. These methods do not rely on gradient information and instead explore the search space based on the function values[jones2020scipy].
@@ 30.4.1 Nelder-Mead Method
The Nelder-Mead method is a popular gradient-free optimization algorithm. It uses a simplex of $n+1$ points in $n$-dimensional space and iteratively adjusts the simplex to converge to the minimum of the function[lee2023nelder]. The algorithm involves operations like reflection, expansion, contraction, and shrinking to adjust the simplex[rios2023derivative].
@@ 30.5 Applications in Training Deep Neural Networks
SGD is a variant of gradient descent that updates the parameters using a small subset (mini-batch) of the training data at each step[robbins1951stochastic, bottou2010large]. This makes it more efficient for large datasets. The update rule for SGD is:
$x_{t+1}=x_{t}-\eta\nabla f(x_{t};\text{mini-batch})$
@@
6class SimpleNet($n$.Module):
@@
Ordinary Differential Equations (ODEs) are equations that describe the relationship between a function and its derivatives[arnold1992ordinary, brown2023ode]. They play a crucial role in many fields of science and engineering, including neural networks and deep learning, where they are used to model dynamic systems[chen2023neural, rossi2024adaptive]. In this chapter, we will cover the basic concepts of ODEs, methods for solving them, and their applications in modeling neural dynamics.
@@ 31.2 Euler’s Method
Euler’s method is the simplest numerical technique for solving ODEs[kincaid2009numerical]. It approximates the solution of an ODE by taking small steps along the curve of the solution. Given the ODE:
@@
  • We define the function f(t,y)=2yf(t,y)=-2y and solve it using Euler’s method.
@@ 31.3 Runge-Kutta Methods
Euler’s method, while simple, is not very accurate for small step sizes. Runge-Kutta methods are more advanced numerical techniques that provide better accuracy for solving ODEs[runge1895ver, dowling2023review]. The most commonly used method is the fourth-order Runge-Kutta method (RK4), which provides a much better approximation of the solution by using a weighted average of several slopes[kutta1901beitrag, martinez2023time, johnson2023rkmethods].
@@ 31.4 Stiff ODEs
Stiff ODEs are a special class of differential equations where certain numerical methods (such as Euler’s method) can become unstable unless very small step sizes are used[gear1971first, hairer1996solving, hairer2009solving]. Stiffness typically arises in systems where there are processes that occur on vastly different time scales[fayed2023new, yang2023adaptive, chen2023dynamical, mccall2024efficient, torres2024review].